OpenStudy (anonymous):

Can someone help me with the first problem on the nuggets code? Having trouble on where to start with problem 1?

OpenStudy (anonymous):

Well, the problem is how to calculate nuggets, from different combinations of the indicated Pack Sizes. I would start by writing down the equation, then write in steps what you would have to do to solve that equation. ex. 1. I need to get total N(nuggets), using equation: 6a + 9b + 20c. 2. I must calculate for N =50, to 55 nuggets inclusive. 3. If I get a solution for N = 50 - 55, I must explain why N = 56 to 65 is also possible. 4. How do I substitute int for those variables (a,b,c)? 5. What Range of values would I need, if any? This is just a basic way that I solved the problem, but there are many other ways to do it.

OpenStudy (anonymous):

Thank you so much I just had no idea how to start it and now I actually have a clue.